OpenStudy (anonymous):

a local vendor is giving away treats. Two baskets contain 12 bags of chips and 10 candy bars. if you randomly select one treat from each basket what is the probability that 1 treat is a bag of chips and 1 treat is a candy bar?

OpenStudy (anonymous):

prob of selecting 1 bag of chip is 1/12 and that of selecting 1 candy bar is 1/10 hence reqd. compound prob = 1/2 * 1/10 =1/120

OpenStudy (anonymous):

total 22c2=231

OpenStudy (anonymous):

i agree with @matricked.

OpenStudy (anonymous):

nope its 120/231

OpenStudy (anonymous):

@kropot72 why u delete that yours is correct

OpenStudy (anonymous):

12c1*10c1/22c2=120/231

OpenStudy (anonymous):

oh um ok.

OpenStudy (anonymous):

got the answer right?

OpenStudy (anonymous):

wait you are talking to me?